# Dynamic angular synchronization under smoothness constraints

### v2

@@ 1 Introduction
In the most typical setting, the group elements $g^{*}_{1},\ldots,g^{*}_{n}$ are represented as the nodes of a simple undirected graph $\mathcal{H}=([n],\mathcal{E})$, where an edge $\{i,j\}$ is present if and only if the observer has access to a noisy version of $g^{*}_{i}{g^{*}_{j}}^{-1}$. In the particular case where $\mathcal{H}$ is connected and the exact value $g^{*}_{i}{g^{*}_{j}}^{-1}$ is observed, the recovery problem has a straightforward solution. Indeed, it suffices to consider a spanning tree of $\mathcal{H}$, assign any value to an arbitrarily chosen root, and the remaining $n-1$ values are found by traversing the edges of tree, solving each time a simple equation. Consequently, these $n-1$ values are uniquely determined up to an global “shift” defined by the chosen value of the root node. On the other hand, if $\mathcal{H}$ is disconnected, there is insufficient information to determine a global solution, since there are no comparisons between the elements of different connected components.
@@ 1.1 Angular synchronization
In the angular synchronization problem, which is the focus of this paper, the goal is to recover $n$ unknown angles $\theta_{1}^{*},\ldots,\theta_{n}^{*}\in[0,2\pi)$ from a set of $m$ noisy measurements of their pairwise offsets $\Theta_{ij}=(\theta_{i}^{*}-\theta_{j}^{*})\mod 2\pi$. Equivalently, we can represent each angle $\theta^{*}_{i}$ by a unit modulus complex number $g^{*}_{i}=e^{\iota\theta^{*}_{i}}$ and the pairwise offsets by $g^{*}_{i}{g^{*}_{j}}^{\mathsf{H}}$, where for any $z\in\mathbb{C}$ , ${z}^{\mathsf{H}}$ represent its complex conjugate. As explained above, the available pairwise measurements can be encoded into the edges of a measurement graph $\mathcal{H}$, which in most typical applications can be modeled as a sparse random geometric graph (e.g., disc graph), a $k$-nearest-neighbor graph, or an Erdős-Rényi graph. The difficulty of the problem is amplified, on one hand, by the amount of noise in the available pairwise offset measurements, and on the other hand by the sparsity of the measurements, since typically, ||($|{\mathcal{E}}|\ll{n\choose 2}$, i.e., only a small subset of all possible pairwise offsets are available to the user.
@@
  • Additive Gaussian noise (AGN) model [5]. Here, $\mathcal{H}$ is the simple complete graph and the observations are of the form $g^{*}_{i}{g^{*}_{j}}^{\mathsf{H}}+\sigma W_{ij}$, where $(W_{ij})_{i<j}$ are independent complex standard Gaussian’s, $W_{ii}=0$ (for all $i$), and $\sigma>0$ represents the noise level. Note that (1.1) is the maximum likelihood estimator (MLE) for the AGN model.
@@ 2.1 Notation
For any $n\in\mathbb{N}$, define $[n]=\{1,\cdots,n\}$. For $A\in\mathbb{C}^{n\times p}$, ${A}^{\mathsf{H}}$ denotes its Hermitian conjugate transpose. We define the operators and Im, from $\mathbb{C}^{n\times p}$ to $\mathbb{R}^{n\times p}$, which for any matrix give their real and imaginary parts, respectively. Denote $e_{1},\dots,e_{n}\in\mathbb{R}^{n}$ to be the canonical vectors. We represent the vector of all ones as $\mathbbm{1}$, with the dimension being clear from the context.
@@ 2.2 Problem setup
$A(k)=(G^{*}(k)-I_{n})+\sigma W(k)\enspace\text{ for }k\in[T],$ (2.2)
@@
$A_{ij}(k)=\begin{cases}g^{*}_{i}(k){g^{*}_{j}(k)}^{\mathsf{H}}\text{ with probability }\left(1-\eta\right)p(k),\\ e^{\iota\varphi_{ij}(k)}\enskip\enskip\quad\text{ with probability }\eta p(k),\\ 0\enskip\qquad\qquad\text{ with probability }(1-p(k)),\end{cases}$
@@
Denoting $M\in\left\{{0,1,-1}\right\}^{T\times(T-1)}$ to be the incidence matrix of any directed path graph [^2] The rows of $M$ correspond to the $T$ vertices, and the columns correspond to the $T-1$ directed edges. For the column corresponding to edge $i\rightarrow j$, we set its $i$th entry to $+1$, the $j$th entry to $-1$, and the rest of the entries in that column to 00. on $T$ vertices, this can be written equivalently as
@@ 2.3.1 Global sphere relaxation
As mentioned in [5, Sec.2], the problem (2.14) is an instance of complex quadratic programming, which is known to be NP-Hard [45, Prop.3.5]. Without the penalty term, problem (2.11) amounts to solve $T$ independent instances of (2.14), clearly remaining NP-Hard. However, it is unclear whether the hardness holds for all $\lambda>0$. In particular, if one allows the value $\lambda=\infty$, the problem becomes trivial as any perfectly smooth $G$ (that is $G(k)=G(1)$ for all $k\in[T]$) will be a solution. Nevertheless, the problem remains highly non-convex even for large values of $\lambda$, which makes the existence of an efficient algorithm for solving (2.11) (for an arbitrary $A$) unlikely.
@@
[…]
@@ 2.3.2 Local sphere relaxation
Algorithm 2 Local TRS $+$ Global Smoothing for Dynamic Synchronization (LTRS-GS-DynSync)
@@ 2.3.3 Denoising measurement matrices
Algorithm 3 Global Matrix Denoising $+$ Local TRS for Dynamic Synchronization (GMD-LTRS-DynSync)
@@ 3.1 Matrix denoising under AGN model
When $\tau=T$, we have $\mathcal{P}^{\perp}_{\tau,n}=0$, from which the final bound follows. So assume now that $11\leq\tau\leq T-$ 1. Denote the $j$-th column of $G^{*}$ by $G^{*}_{:,j}$. Since
@@
This implies (notice that $\lambda_{T-\tau}\neq 0$ since $11\leq\tau\leq T-$ 1)
@@
Observe that $W^{\prime}_{1}$ has independent entries, with each entry either 00 or drawn from $\mathcal{N}(0,1/2)$. To bound $\mathcal{V}_{1}$, we will use the Hanson-Wright inequality [39, Thm. 1.1]. It is easy to see that
@@
Finally, we remark that the $O(T^{2/3}S_{T}^{1/3})$ term in (3.7) matches that obtained in [3, Theorem B] and also matches the rate for denoising smooth signals on a path graph [40, Theorem 6].
@@ 3.2 Matrix denoising under the Outliers model
$\|\mathcal{P}_{\tau,n}(DG^{*})\|_{F}^{2}\leq\mu\bar{d}^{2}\|\mathcal{P}_{\tau,n}G^{*}\|_{F}^{2},$
@@
$\|\mathcal{P}_{\tau,n}(DG^{*})\|_{F}^{2}\leq C_{1}\frac{\mu\bar{d}^{2}nT^{2}S_{T}}{\tau^{2}}\mathbbm{1}_{\tau<T}$ (3.11)
@@
Moreover, $Q(\cdot)$ is non-negative continuous on $[0,1]$, with $Q(0+0)=Q(1-0)=0$, and $Q^{2}(1/2)=1/8$ (see [37]). It is easy to also verify that $Q(p)\leq\frac{1}{\sqrt{2\log(1/p)}}$ if $p\leq 1/4$.
@@ 3.3 Spectral local synchronization guarantee
Algorithm 4 Global Matrix Denoising $+$ Local Spectral method for Dynamic Synchronization (GMD-LSPEC-DynSync)
@@
    In particular, when $\sigma>0$ is a constant, then $\|\widehat{g}-g^{*}\|_{2}^{2}=O(T^{2/3}S_{T}^{1/3})$ which matches the (optimal) rate for denoising smooth signals on a path graph on $T$ vertices [40, Theorems 5,6]; recall the discussion in Remark 4.
@@
The bound exhibits the same scaling [^4] Note that the rates in [40, Theorems 5,6] are obtained for additive Gaussian noise. We believe the dependency on $T$ should be unchanged for more general sub-Gaussian noise models, however this needs to be established formally. w.r.t $T,S_{T}$ as in the AGN model; recall Remark 11.
@@ 4 Experiments
    In order to choose the appropriate value of the regularization parameters for the algorithms ($\lambda$ for GTRS-DynSyncand $\tau$ for LTRS-GS-DynSync, GMD-LTRS-DynSync), we consider a data-driven grid-search based heuristic. In particular, define a discretization of $[0,T]$, where (i) the first $\lfloor\sqrt{T}\rfloor+1$ elements of the grid are the numbers $0,1,\dots,\lfloor\sqrt{T}\rfloor$; (ii) the next five elements are equi-spaced numbers in the interval $[\lfloor\sqrt{T}\rfloor,\lfloor T^{2/3}\rfloor]$; (iii) the last five elements are equi-spaced numbers in the interval $[\lfloor T^{2/3}\rfloor,T]$. For each value on the grid, call it $\beta$, we set $\tau=\min\left\{{\beta+1,T}\right\}$ and $\lambda=\beta\times\lambda_{\text{scale}}$ for a fixed choice of the scaling parameter $\lambda_{\text{scale}}>0$. For the corresponding output $\widehat{g}$ of each algorithm, we compute a ‘data-fidelity’ value using the input pairwise data, as for the data-fidelity term in (2.13) (with $\tilde{g}$ therein replaced by $\widehat{g}$). Then we select the optimal $\beta$ (thus optimal $\tau$ and $\lambda$) for each algorithm as follows.
@@
Figure 3: RMSE versus $T$ for AGN model ($n=30$, $\sigma=3$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over $20$ MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@
Figure 4: RMSE versus $T$ for Outliers model ($n=30$, $\eta=p=0.2$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over $20$ MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@
In this experiment, we evaluate the RMSE versus the noise-level for all the above methods. The noise-level is denoted by $\gamma$ for ease of exposition, with $\gamma=\sigma>0$ for the AGN model (varied from 00 to $10$), and $\gamma=\eta\in[0,1]$ for the Outliers model (varied from 00 to $0.8$). Throughout, we fix $n=30$, $T=20$, and perform $20$ Monte Carlo (MC) runs for each value of $\gamma$. In a single MC run, we randomly generate the ground-truth for a specified smoothness $S_{T}$ (as described earlier), then generate the noisy pairwise data (as per specified noise model), and then compute the RMSE for each algorithm for the regularization parameters selected by the aforementioned data-fidelity rule. The RMSE values are then averaged over the $20$ runs. The results are shown in Fig. 5 for the Wigner model, and Fig. 6 for the Outliers model (with $p=0.2$). The following observations can be made from these plots.
@@
    For the AGN model (Fig. 5), note that the performance of the spectral method and GTRS-DynSync are similar for low noise levels ($\sigma$ less than $1$), both being worse than the other two methods in this noise regime. The performance of the spectral method degrades for $\sigma>1$, while that of the other three methods is similar at moderate noise levels ($51<\sigma<$ 5). For larger values of $\sigma$, we see that GMD-LTRS-DynSync performs best, while LTRS-GS-DynSync and GTRS-DynSync have similar error values.
@@
Figure 5: RMSE versus $\gamma(=\sigma)$ for AGN model ($n=30$, $T=20$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over $20$ MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@
Figure 6: RMSE versus $\gamma(=\eta)$ for Outliers model ($n=30$, $p=0.2$, $T=20$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over $20$ MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.

### v3

@@ 1 Introduction
In the most typical setting, the group elements $g^{*}_{1},\ldots,g^{*}_{n}$ are represented as the nodes of a simple undirected graph $\mathcal{H}=([n],\mathcal{E})$, where an edge $\{i,j\}$ is present if and only if the observer has access to a noisy version of $g^{*}_{i}{g^{*}_{j}}^{-1}$. In the particular case where $\mathcal{H}$ is connected and the exact value $g^{*}_{i}{g^{*}_{j}}^{-1}$ is observed, the recovery problem has a straightforward solution. Indeed, it suffices to consider a spanning tree of $\mathcal{H}$, assign any value to an arbitrarily chosen root, and the remaining n1n-1 values are found by traversing the edges of tree, solving each time a simple equation. Consequently, these n1n-1 values are uniquely determined up to an global “shift” defined by the chosen value of the root node. On the other hand, if $\mathcal{H}$ is disconnected, there is insufficient information to determine a global solution, since there are no comparisons between the elements of different connected components.
@@ 1.1 Angular synchronization
In the angular synchronization problem, which is the focus of this paper, the goal is to recover $n$ unknown angles $\theta_{1}^{*},\ldots,\theta_{n}^{*}\in[0,2\pi)$ from a set of $m$ noisy measurements of their pairwise offsets $\Theta_{ij}=(\theta_{i}^{*}-\theta_{j}^{*})\mod 2\pi$. Equivalently, we can represent each angle $\theta^{*}_{i}$ by a unit modulus complex number $g^{*}_{i}=e^{\iota\theta^{*}_{i}}$ and the pairwise offsets by $g^{*}_{i}{g^{*}_{j}}^{\mathsf{H}}$, where for any $z\in\mathbb{C}$ , ${z}^{\mathsf{H}}$ represent its complex conjugate. As explained above, the available pairwise measurements can be encoded into the edges of a measurement graph $\mathcal{H}$, which in most typical applications can be modeled as a sparse random geometric graph (e.g., disc graph), a $k$-nearest-neighbor graph, or an Erdős-Rényi graph. The difficulty of the problem is amplified, on one hand, by the amount of noise in the available pairwise offset measurements, and on the other hand by the sparsity of the measurements, since typically, $|{\mathcal{E}}|\ll{n\choose 2}$, i.e., only a small subset of all possible pairwise offsets are available to the user.
@@
  • Additive Gaussian noise (AGN) model [5]. Here, $\mathcal{H}$ is the simple complete graph and the observations are of the form $g^{*}_{i}{g^{*}_{j}}^{\mathsf{H}}+\sigma W_{ij}$, where $(W_{ij})_{i<j}$ are independent complex standard Gaussian’s, $W_{ii}=0$ (for all ii), and $\sigma>0$ represents the noise level. Note that (1.1) is the maximum likelihood estimator (MLE) for the AGN model.
@@ 2.1 Notation
For any $n\in\mathbb{N}$, define $n]=\{1,\cdots,n\}$. For $A\in\mathbb{C}^{n\times p}$, ${A}^{\mathsf{H}}$ denotes its Hermitian conjugate transpose. We define the operators and Im, from $\mathbb{C}^{n\times p}$ to $\mathbb{R}^{n\times p}$, which for any matrix give their real and imaginary parts, respectively. Denote $e_{1},\dots,e_{n}\in\mathbb{R}^{n}$ to be the canonical vectors. We represent the vector of all ones as $\mathbbm{1}$, with the dimension being clear from the context.
@@ 2.2 Problem setup
$A(k)=(G^{*}(k)-I_{n})+\sigma W(k)\kern 5.0pt\text{ for }k\in[T],$ (2.2)
@@
[…]
@@
Denoting $M\in\left\{{0,1,-1}\right\}^{T\times(T-1)}$ to be the incidence matrix of any directed path graph [^2] The rows of $M$ correspond to the $T$ vertices, and the columns correspond to the T1T-1 directed edges. For the column corresponding to edge $i\rightarrow j$, we set its $i$th entry to +1+1, the $j$th entry to 1-1, and the rest of the entries in that column to 0. on $T$ vertices, this can be written equivalently as
@@ 2.3.1 Global sphere relaxation
As mentioned in [5, Sec.2], the problem (2.14) is an instance of complex quadratic programming, which is known to be NP-Hard [46, Prop.3.5]. Without the penalty term, problem (2.11) amounts to solve $T$ independent instances of (2.14), clearly remaining NP-Hard. However, it is unclear whether the hardness holds for all $\lambda>0$. In particular, if one allows the value $\lambda=\infty$, the problem becomes trivial as any perfectly smooth $G$ (that is $G(k)=G(1)$ for all $k\in[T]$) will be a solution. Nevertheless, the problem remains highly non-convex even for large values of $\lambda$, which makes the existence of an efficient algorithm for solving (2.11) (for an arbitrary $A$) unlikely.
@@
3:Obtain $\widehat{\tilde{g}}=\operatorname{concat}(\widehat{\tilde{g}}(1),\cdots,\widehat{\tilde{g}}(T))\in\mathbb{C}^{(n-1)T}\leftarrow$ a solution of (2.15)
@@ 2.3.2 Local sphere relaxation
Algorithm 2 Local TRS ++ Global Smoothing for Dynamic Synchronization (LTRS-GS-DynSync)
@@ 2.3.3 Denoising measurement matrices
Algorithm 3 Global Matrix Denoising ++ Local TRS for Dynamic Synchronization (GMD-LTRS-DynSync)
@@ 3.1 Matrix denoising under AGN model
When $\tau=T$, we have $\mathcal{P}^{\perp}_{\tau,n}=0$, from which the final bound follows. So assume now that $1\leq\tau\leq T-1$. Denote the $j$-th column of $G^{*}$ by $G^{*}_{:,j}$. Since
@@
This implies (notice that $\lambda_{T-\tau}\neq 0$ since $1\leq\tau\leq T-1$)
@@
Observe that $W^{\prime}_{1}$ has independent entries, with each entry either 0 or drawn from $\mathcal{N}(0,1/2)$. To bound $\mathcal{V}_{1}$, we will use the Hanson-Wright inequality [40, Thm. 1.1]. It is easy to see that
@@
Finally, we remark that the $O(T^{2/3}S_{T}^{1/3})$ term in (3.7) matches that obtained in [3, Theorem B] and also matches the rate for denoising smooth signals on a path graph [41, Theorem 6].
@@ 3.2 Matrix denoising under the Outliers model
$\|{\color[rgb]{0,0,0}\mathcal{P}^{\perp}_{\tau,n}(D\overline{G}^{*})}\|_{F}^{2}\leq\mu\bar{d}^{2}\|{\color[rgb]{0,0,0}\mathcal{P}^{\perp}_{\tau,n}\overline{G}^{*}}\|_{F}^{2},$
@@
$\|{\color[rgb]{0,0,0}\mathcal{P}^{\perp}_{\tau,n}(D\overline{G}^{*})}\|_{F}^{2}\leq C_{1}\frac{\mu\bar{d}^{2}nT^{2}S_{T}}{\tau^{2}}\mathbbm{1}_{\tau<T}$ (3.11)
@@
Moreover, $Q(\cdot)$ is non-negative continuous on $[0,1]$, with Q(0+0)=Q(10)=0Q(0+0)=Q(1-0)=0, and $Q^{2}(1/2)=1/8$ (see [37]). It is easy to also verify that $Q(p)\leq\frac{1}{\sqrt{2\log(1/p)}}$ if $p\leq 1/4$.
@@ 3.3 Spectral local synchronization guarantee
Algorithm 4 Global Matrix Denoising ++ Local Spectral method for Dynamic Synchronization (GMD-LSPEC-DynSync)
@@
    In particular, when $\sigma>0$ is a constant, then $\|\widehat{g}-g^{*}\|_{2}^{2}=O(T^{2/3}S_{T}^{1/3})$ which matches the (optimal) rate for denoising smooth signals on a path graph on $T$ vertices [41, Theorems 5,6]; recall the discussion in Remark 4.
@@
The bound exhibits the same scaling [^4] Note that the rates in [41, Theorems 5,6] are obtained for additive Gaussian noise. We believe the dependency on $T$ should be unchanged for more general sub-Gaussian noise models, however this needs to be established formally. w.r.t $T,S_{T}$ as in the AGN model; recall Remark 11.
@@ 4 Experiments
    In order to choose the appropriate value of the regularization parameters for the algorithms ($\lambda$ for GTRS-DynSyncand $\tau$ for LTRS-GS-DynSync, GMD-LTRS-DynSync), we consider a data-driven grid-search based heuristic. In particular, define a discretization of $[0,T]$, where (i) the first $\lfloor\sqrt{T}\rfloor+1$ elements of the grid are the numbers $0,1,\dots,\lfloor\sqrt{T}\rfloor$; (ii) the next five elements are equi-spaced numbers in the interval $\lfloor\sqrt{T}\rfloor,\lfloor T^{2/3}\rfloor]$; (iii) the last five elements are equi-spaced numbers in the interval $\lfloor T^{2/3}\rfloor,T]$. For each value on the grid, call it $\beta$, we set $\tau=\min\left\{{\beta+1,T}\right\}$ and $\lambda=\beta\times\lambda_{\text{scale}}$ for a fixed choice of the scaling parameter $\lambda_{\text{scale}}>0$. For the corresponding output $\widehat{g}$ of each algorithm, we compute a ‘data-fidelity’ value using the input pairwise data, as for the data-fidelity term in (2.13) (with $\tilde{g}$ therein replaced by $\widehat{g}$). Then we select the optimal $\beta$ (thus optimal $\tau$ and $\lambda$) for each algorithm as follows.
@@
Figure 3: RMSE versus $T$ for AGN model ($n=30$, $\sigma=3$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over 2020 MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@
Figure 4: RMSE versus $T$ for Outliers model ($n=30$, $\eta=p=0.2$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over 2020 MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@
In this experiment, we evaluate the RMSE versus the noise-level for all the above methods. The noise-level is denoted by $\gamma$ for ease of exposition, with $\gamma=\sigma>0$ for the AGN model (varied from 0 to $10$), and $\gamma=\eta\in[0,1]$ for the Outliers model (varied from 0 to $0.8$). Throughout, we fix $n=30$, $T=20$, and perform $20$ Monte Carlo (MC) runs for each value of $\gamma$. In a single MC run, we randomly generate the ground-truth for a specified smoothness $S_{T}$ (as described earlier), then generate the noisy pairwise data (as per specified noise model), and then compute the RMSE for each algorithm for the regularization parameters selected by the aforementioned data-fidelity rule. The RMSE values are then averaged over the $20$ runs. The results are shown in Fig. 5 for the Wigner model, and Fig. 6 for the Outliers model (with $p=0.2$). The following observations can be made from these plots.
@@
    For the AGN model (Fig. 5), note that the performance of the spectral method and GTRS-DynSync are similar for low noise levels ($\sigma$ less than $1$), both being worse than the other two methods in this noise regime. The performance of the spectral method degrades for $\sigma>1$, while that of the other three methods is similar at moderate noise levels ($1<\sigma<5$). For larger values of $\sigma$, we see that GMD-LTRS-DynSync performs best, while LTRS-GS-DynSync and GTRS-DynSync have similar error values.
@@
Figure 5: RMSE versus $\gamma(=\sigma)$ for AGN model ($n=30$, $T=20$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over 2020 MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@
Figure 6: RMSE versus $\gamma(=\eta)$ for Outliers model ($n=30$, $p=0.2$, $T=20$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over 2020 MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.